# FLASC: A Flare-Sensitive Clustering Algorithm

### v1

@@ 3.1. The FLASC Algorithm
5:         compute the centrality $c(\mathbf{x}_{i})$ for all $\mathbf{x}_{i}\in Cj$.
6:         extract the cluster centrality graph $\mathbf{G}_{c}^{C_{j}}$.
7:         compute the Single Linkage Clustering hierarchy of $\mathbf{G}_{c}^{C_{j}}$
@@
Detecting flares in a cluster centrality graph $\mathbf{G}_{c}^{C_{j}}$ is conceptually similar to detecting density-based clusters. Instead of the density profile (as in Figure 2), we now look at the eccentricity profile ($e(\mathbf{x}_{i})=1/c(\mathbf{x}_{i})$). The contour clusters in this profile correspond to the flares we want to detect. Effectively, we perform Single Linkage Clustering (Sibson, 1973) on the cluster centrality graph $\mathbf{G}_{c}^{C_{j}}$ using a Union-Find data structure as in (McInnes and Healy, 2017). The resulting hierarchy is simplified using a minimum branch size $m_{b}$.
@@
HDBSCAN*’s ‘flat’ clustering strategies are used to compute branch labels and membership probabilities from these condensed hierarchies. Points that enter the filtration after the selected flares have connected—i.e. points with the noise label—are given a single non-noise label representing the cluster’s centre. Finally, the cluster and branch labels are combined (see Figure 1). By default, points in clusters with two or fewer branches are given a single label because two branches are expected in all clusters, indicating the outsides growing towards each other. The label sides as branches parameter can be used to turn off this behaviour and separate the ends of elongated clusters in the labelling. The cluster and branch probabilities are combined by taking their average value (see Figure 4).
@@
Other labelling and probability combinations are possible. For example, the cluster and branch probability product more strongly emphasises the outsides of the flares (see Figure 4). Like McInnes et al.’s HDBSCAN* implementation (McInnes et al., 2017), FLASC supports computing flare membership vectors, describing how much a point $\mathbf{x}_{i}\in C_{j}$ belongs to each branch $B_{b}\subset C_{j}$. These membership values are based on the geodesic distances in the cluster approximation graph $\mathbf{G}_{k}^{C_{j}}$: $d_{geo}(\mathbf{r}_{B_{b}},\mathbf{x}_{i})$, where $\mathbf{r}_{B_{b}}$ is the branch’ root, i.e., the point closest to the branch’s weighted average $\bar{\mathbf{x}}_{B_{b}}$, where the probability assigned to each point is used as weight. The flare membership vectors can be used to label points by the closest branch root, as in Figure 4, which can change the label for central points. Alternatively, a softmax function can be used to convert $d_{geo}(\mathbf{r}_{B_{b}},\mathbf{x}_{i})$ into the membership probabilities:
(4) $p(\mathbf{x}_{i},B_{b})=\frac{e^{c_{b}(\mathbf{x}_{i},B_{b})}}{\sum_{B_{l}\in C_{j}}{e^{c_{b}(\mathbf{x}_{i},B_{l})}}},$
where $c_{b}(\mathbf{x}_{i},B_{b})$ converts $d_{geo}(\mathbf{r}_{B_{b}},\mathbf{x}_{i})$ into a flare centrality as in (3) (see Figure 4).
@@
Low persistent branches can be ignored using a branch selection persistence parameter, analogous to HDBSCAN*’s cluster selection epsilon (Malzer and Baum, 2020). However, flares do not have to start at zero centrality, so branch selection persistence describes the minimum centrality range rather than a single centrality threshold value. The procedure for applying this persistence threshold evaluates selected branches in increasing order of persistence to avoid parent-child relations in the final chosen branches. In addition, the flare membership probability is based on the flare’s maximum eccentricity rather than the selected segment’s maximum eccentricity in the branch condensed tree, as is used for clusters.
@@
Figure 4. Different ways to combine cluster and branch membership probabilities. The cluster and branch probability average (a) and product (b) are visualised with desaturation. (c) Points labelled by the geodesically closest branch root—i.e., the point closest to the branch’s weighted average—and desaturated as in (a). (d) Weighted branch membership for the orange branch is visualised by transparency. Branch memberships are computed from the traversal distance to the branch’s root.
Four scatterplots of a Y-shaped cluster with different cluster and branch labelling and probability combinations.
Figure 4. Different ways to combine cluster and branch membership probabilities. The cluster and branch probability average (a) and product (b) are visualised with desaturation. (c) Points labelled by the geodesically closest branch root—i.e., the point closest to the branch’s weighted average—and desaturated as in (a). (d) Weighted branch membership for the orange branch is visualised by transparency. Branch memberships are computed from the traversal distance to the branch’s root.
3.2. Stability
@@
Two notions of stability are relevant: (1) the algorithm has to provide similar results when run repeatedly on (different) samples of an underlying distribution, and (2) the detected branch hierarchies have to represent the clusters’ underlying topology accurately.
@@
Vandaele et al. analysed the latter notion of stability for graph-based branch detection, explaining that the graph approximation should accurately represent the underlying shape and the graph-based centrality function should accurately describe the points’ centrality in a cluster’s metric space $(C_{j},d_{mreach})$. For the normalised centrality used by Carlsson (Carlsson, 2014), they showed that the bound on the bottleneck distances between true and empirical persistence diagrams is tight if the metric distortion induced by the graph and its maximum edge weight is small.
@@
The centrality metric (Equation (3)) is more complex to analyse. It is a $\mathrm{c}$-Lipschitz-continuous function when considered over the cluster centrality graph’s edges:
$|\max{\{c(\mathbf{x}_{i}),c(\mathbf{x}_{j})\}}-\max{\{c(\mathbf{x}_{k}),c(\mathbf{x}_{l})\}}|\leq\mathrm{c}\;d_{mreach}(\mathbf{x}_{i},\mathbf{x}_{l}),$
@@
where $\mathrm{c}$ is a constant describing the continuity, $(\mathbf{x}_{i},\mathbf{x}_{j})\in\mathbf{G}_{c}^{C_{j}}$ and $(\mathbf{x}_{k},\mathbf{x}_{l})\in\mathbf{G}_{c}^{C_{j}}$, and the mutual reachability between $\mathbf{x}_{i}$ and $\mathbf{x}_{l}$ is the largest of the four points. However, the position of the cluster’s centroid $\bar{\mathbf{x}}_{C_{j}}$ influences how well the function represents the centrality in the cluster’s metric space $(C_{j},d_{mreach})$. We aim to show the current approach strikes a good balance between computational cost and stability in the experiments presented in Section 4.
3.3. Complexity
@@
The algorithm’s most computationally expensive steps are constructing the full cluster approximation graphs and computing the cluster centrality graphs’ single linkage hierarchies. Naively, the worst-case complexity for creating a cluster approximation graph is $\mathcal{O}(n_{c}^{2})$, where $n_{c}$ is the number of points in the cluster. Usually, the average case is much better because the approximation graphs rarely are fully connected. After all, HDBSCAN*’s noise classification limits the density range within the clusters. Furthermore, the space tree that we re-use from the HDBSCAN* clustering step provides fast asymptotic performance for finding the graph’s edges. The exact run-time bounds depend on the data properties. They are challenging to describe (as explained in (McInnes and Healy, 2017)), but an average complexity proportional to $\mathcal{O}(n_{e}\log{N})$ is expected, where $n_{e}$ is the number of edges in the graph. Computing single linkage hierarchies from the cluster centrality graphs is possible in $\mathcal{O}(n_{e}\alpha(n_{e}))$ using McInnes et al. (McInnes et al., 2017)’s Union-Find implementation adapted to ignore edges between data points that are already in the same connected component. Like McInnes and Healy (McInnes and Healy, 2017), we feel confident that FLASC achieves sub-quadratic complexity on average, which we demonstrate in a practical example in Section 4.4.
4. Experiments and Results
This section presents two case studies demonstrating how FLASC can be used on real-world data to detect branch-based subgroups. In addition, we describe two analyses of synthetic data sets to demonstrate FLASC’s stability and computational performance.
4.1. Case Study: Diabetes Types
In 1979, Reaven and Miller (Reaven and Miller, 1979) collected and analysed a data set about diabetes patients. Previously, they had found a “horse shoe”-shaped relation between glucose levels and insulin responses in diabetes patients (Reaven and Miller, 1979). Their 1979 paper attempted to clarify that relationship by measuring additional metabolic variables. Three of those variables turned out to be very informative in a 3D scatterplot. Our recreation of that scatterplot is shown in Figure 5(a), illustrating a dense core with two less-dense branches, which Reaven and Miller considered unlikely to be a single population (Reaven and Miller, 1979).
More recently, Singh et al. used the data set to demonstrate how Mapper with a density-based lens function visualises these flares without manually specifying which dimensions to plot (Singh et al., 2007). Their analysis leverages the flares’ lower density, allowing them to be detected without a centrality metric. In general, though, local density minima do not always relate to branches, especially for data sets with multiple branching clusters.
In this section, we show how FLASC can detect the branching pattern in this data set and classify the observations by their branch without manually extracting the flares from a visualisation.
4.1.1. Evaluation and settings
The data set—obtained from (Andrews and Herzberg, 1985)—contains five variables describing 145 subjects: the relative weight, the plasma glucose level after a period of fasting, the steady-state plasma glucose response (SSPG), and two areas under a curve—one for glucose (AUGC) and one for insulin (AUIC)—representing the total amount observed during the experimental procedure described in (Reaven and Miller, 1979). All five variables were z-score normalised and used to compute the Euclidean distance between subjects.
4.1.2. Results
Figure 5. (a) A three-dimensional scatterplot showing the area under the plasma glucose curve (AUGC), the area under the insulin curve (AUIC), and the steady state plasma glucose response (SSPG) as described by (Reaven and Miller, 1979). (b) FLASC’s classification tuned to find one cluster with multiple branches. One low-persistent branch (blue) corresponding to the bottom of the centre could be filtered out by specifying a persistence threshold. (c) HDBSCAN*’s labels tuned to find multiple clusters. Grey points were classified as noise.
(a) Illustration
(b) FLASC
(c) HDBSCAN*
Three variants of the 3D scatterplot indicating how the subgroups that HDBSCAN* and FLASC detect relate to Reaven and Millers interpretation.
Figure 5. (a) A three-dimensional scatterplot showing the area under the plasma glucose curve (AUGC), the area under the insulin curve (AUIC), and the steady state plasma glucose response (SSPG) as described by (Reaven and Miller, 1979). (b) FLASC’s classification tuned to find one cluster with multiple branches. One low-persistent branch (blue) corresponding to the bottom of the centre could be filtered out by specifying a persistence threshold. (c) HDBSCAN*’s labels tuned to find multiple clusters. Grey points were classified as noise.
Figure 5 shows the resulting classifications encoded as colour on the 3D scatterplot. FLASC’s classification (Figure 5(b)) distinguishes the flares from the central core. The algorithm also finds a low-persistent flare representing the central core’s bottom. This flare could be ignored by specifying a persistence threshold. In contrast, as expected, HDBSCAN*’s classification (Figure 5(c)) does not find the flares. Instead, it finds part of the left flare as a small low-persistent cluster and merges most of the right flare with the central core. This demonstrates that branching structures cannot be detected as density-based clusters when they do not contain local density maxima. In this case, though, specifying a larger minimum cluster size and re-clustering the points classified as noise should detect the two branch-based subgroups because these branches have a lower density than the central core.
4.2. Case Study: Cell Development
The small roundworm C. Elegans is often used in biological studies and was the first animal to have its genome sequenced completely (elegans Sequencing Consortium*, 1998). More recently, Packer et al. analysed gene expressions in C. Elegans embryos through single-cell RNA sequencing to uncover the trajectories along which cells develop (Packer et al., 2019). Broadly speaking, this data set describes what happens in cells as they develop from a single egg cell into all the different tissues within fully grown C. Elegans worms.
After pre-processing, the data set appears to contain both cluster and branching structures when viewed in Packer et al.’s 3D projection (Packer et al., 2019). In this case study, we demonstrate that FLASC can provide additional information about a data set’s structure even when the main subgroups can be detected as clusters. This information, however, is limited to the most eccentric connections between branches.
4.2.1. Evaluation and settings
The data and pre-processing scripts were obtained from Monocle 3’s (Cao et al., 2019) documentation (Pliner et al., 2022). The pre-processing stages normalise the data, extract the 50 strongest PCA components, and correct for batch effects using algorithms from (Haghverdi et al., 2018). HDBSCAN* and FLASC were evaluated on the pre-processed data using the angular distance metric because neither algorithm supports the cosine distance metric in their optimised code paths. HDBSCAN* was tuned to find multiple clusters with min samples $k$ and min cluster size $m_{c}$ set to $20$. FLASC was tuned to find a single cluster by selecting min samples $k=5$, min cluster size $m_{c}=20$, min branch size $m_{b}=20$, enabling allow single cluster, and setting cluster selection epsilon to $1/1.4$ to classify points as noise if they are ‘too far away’ from the clusters. Branches were detected using the core cluster approximation graph.
4.2.2. Results
Figures 6(a) and 6(b) show the pre-processed data projected into two dimensions using densMAP (Narayan et al., 2021). Data points are coloured to indicate the detected flares and clusters, respectively. In Figure 6(a), the core cluster approximation graph’s edges are also drawn. Notice how some parts of the projection are not connected as one might assume from their coordinates. For example, there are edges between branches 1 and 5 and branches 12 and 14.
Figure 6. 2D densMAP (Narayan et al., 2021) projections of the 50 dimensional pre-processed single-cell gene expression data from (Packer et al., 2019) coloured by the detected branches (a) and clusters (b). (c) and (d) show the branch and cluster condensed trees as icicle plots adapted from (McInnes et al., 2017). The selected branches and clusters are indicated with numbers and coloured. The cluster approximation graph’s edges are drawn on top of the 2d projection in (a).
(a) FLASC
(b) HDBSCAN*
(c) Branch hierarchy
(d) Cluster hierarchy
The clusters that HDBSCAN* and branches that FLASC detect on the single-cell sequencing data. The cluster and branch hierarchies partially communicate the point cloud’s shape.
Figure 6. 2D densMAP (Narayan et al., 2021) projections of the 50 dimensional pre-processed single-cell gene expression data from (Packer et al., 2019) coloured by the detected branches (a) and clusters (b). (c) and (d) show the branch and cluster condensed trees as icicle plots adapted from (McInnes et al., 2017). The selected branches and clusters are indicated with numbers and coloured. The cluster approximation graph’s edges are drawn on top of the 2d projection in (a).
Figures 6(c) and 6(d) visualise the branch and cluster condensed trees, respectively, using an icicle plot adapted from McInnes et al. (McInnes et al., 2017). Segment widths encode the number of points in the tree below the segment. Numbers and colours indicate the selected branches and clusters.
In contrast to the previous case study, HDBSCAN* appears to detect all flares as clusters, indicating that the branches have a local density maximum. Considering that the branches correspond to cell fates—i.e., the developmental end-states—it is unsurprising that local density maxima occur within them. One could imagine that the variation in gene expression is higher during development and that fully developed cells are observed more frequently. Both scenarios could cause these local density maxima. HDBSCAN* also classifies a lot of points as noise. Consequently, it does not capture the trajectories: it only detects the subgroups and does not provide information about how they relate. For example, clusters 5 and 6 are neighbours in the cluster-condensed tree (Figure 6(d)), but the trajectory between those clusters would pass through cluster 16.
FLASC finds the same subgroups, except that clusters 5 and 6 are combined into branch 11 and clusters 13 and 15 into branch 4. In addition, the branch condensed tree (Figure 6(c)) captures (some of) the cluster’s shape. Branches that merge into the cluster near each other are also close in the branch-condensed tree. See, for example, branches 6 and 8. For branches connected to multiple other branches in the cluster approximation graph, only the most eccentric connection is captured by the branch-condensed tree. For example, branches 12 and 14 are neighbours in the hierarchy, while their connections to other branches are more apparent in the densMAP projection (Figure 6(a)).
All in all, this case study demonstrated that FLASC’s branch hierarchy can provide information about a cluster’s shape that may not be obvious in 2D projections.
4.3. Flare Stability
In this section, we explore FLASC’s stability in terms of its difference in output on multiple samples of the same underlying distribution and its accuracy compared to a ground truth.
4.3.1. Datasets
@@
For this comparison, two-dimensional data sets containing a single cluster with $3$ branches laid out as a three-pointed star were generated (Figure 7). The branches span from the centre outwards and are spread out equally across the $2$D plane—i.e., the angle between adjacent branches is $120$ degrees. Each branch has a length $b_{l}$ and contains $100$ points exponentially spaced from the inside out. Consequently, the density is highest at the cluster’s centre and lowest in the branch ends, and this difference increases with the branch length. Normally distributed noise ($\mu=0$) is added to the points’ coordinates using a noise ratio parameter $n_{r}$ to determine the distribution’s standard deviation $\sigma$:
(5) $n_{r}=\frac{4\sigma}{\sqrt{\frac{1}{2}}b_{l}}.$
@@
Here, $n_{r}$ indicates that approximately $95\%$ of the sampled noise values fall within ±$\pm\sqrt{\frac{1}{2}}n_{r}b_{l}$ of zero. Consequently, approximately $90\%$ of points are moved less than $n_{r}b_{l}$ from their original 2D position. By design, the branches do not reliably contain local density maxima, making them practically undetectable as density-based clusters (see Figure 11).
4.3.2. Evaluation and settings
@@
The data sets were sampled with varying branch lengths $b_{l}$ ($2$ to $100$ in $10$ exponentially spaced steps) and noise ratios $n_{r}$ (00 to $1$ in $10$ exponentially spaced steps). Ten data sets were sampled for each parameter combination, resulting in $1000$ point clouds (see Figure 8 for one example of each combination). FLASC was evaluated with full and core approximation graphs on each data set, allow single cluster enabled, min samples $k=5$, min cluster size $m_{c}=25$, a varying min branch size $m_{b}$ ($2$ to $24$ in steps of $2$), and both the eom and leaf branch selection strategies.
@@
For each evaluation, the resulting data point labels and probabilities were stored. From these values, we computed the Adjusted Rand Index (ARI) (Hubert and Arabie, 1985) to describe the agreement between ground truth and assigned partitioning labels adjusted for chance. In addition, for each subgroup detected by FLASC, we computed the weighted average coordinates—i.e., found centroids—to describe the algorithm’s stability. These centroids were assigned to the closest ground truth centroid, and their (unweighted) average position was computed for each ground truth group over the $10$ data sets with the same branch length $b_{l}$ and noise ratio $n_{r}$. The distances between the centroids and their (unweighted) average normalised by the branch length $b_{l}$ serve as a stability metric referred to as the centroid spread (see Figure 7).

### v2

@@ Introduction
5:         compute the eccentricity $e(\mathbf{x}_{i})$ for all $\mathbf{x}_{i}\in Cj$.
6:         extract the cluster approximation graph $\mathbf{G}_{k}^{C_{j}}$.
7:         compute the single linkage clustering hierarchy of $\mathbf{G}_{k}^{C_{j}}$
@@
The main contribution of our work is a flare detection post-processing step for HDBSCAN*. This section describes how the post-processing step works and integrates with HDBSCAN* to form our FLASC algorithm (see Algorithm 1). FLASC starts by evaluating a flat HDBSCAN* clustering, keeping track of the space tree used in HDBSCAN* (McInnes et al., 2017; McInnes and Healy, 2017) to efficiently find nearest neighbours. One noteworthy change from McInnes et al. (2017)’s implementation is that we give all points the 00-label when a single cluster is allowed and selected and no cluster selection epsilon is applied. This enables FLASC to better analyse branching structures in data sets that contain a single cluster. Then, for each selected cluster $C_{j}$, a branch detection step is performed, explained in more detail below.
The concepts behind density-based clustering can also be applied to detect branches within clusters by using an eccentricity measure in place of density, as shown in Fig. 2. Peaks in an eccentricity profile correspond to branches in the cluster, as shown in Fig. 2(b) and Fig. 2(c). We define eccentricity as the distance to the cluster’s centroid:
$e(\mathbf{x}_{i})=d(\bar{\mathbf{x}}_{C_{j}},\mathbf{x}_{i}),$ (3)
where $\bar{\mathbf{x}}_{C_{j}}$ is the cluster’s membership-weighted average (Fig. 2(a)). This eccentricity measure can be computed in $\mathcal{O}(N)$. Comparable to density contour clusters, an eccentricity contour cluster is a maximal, connected subset of points with an eccentricity larger than some threshold $\{\mathbf{x}\;|\;e(\mathbf{x})\geq e_{t}\}$. As in functional persistence (Carlsson, 2014), eccentricity thresholds filter out cluster cores, which separates branches and makes them detectable as connected components.
Similar to HDBSCAN*, we need a notion of connectivity between data points to determine whether two points are part of the same eccentricity contour cluster at a threshold $e_{t}$. In FLASC, we provide two solutions based on the cluster’s density scale in the form of cluster approximation graphs $\mathbf{G}_{k}^{C_{j}}$: the full approximation graph and the core approximation graph. Both types contain a vertex for each point in the cluster $\mathbf{x}_{i}\in C_{j}$ but differ in which edges they include. The full approximation graph adds all edges with $d_{mreach}(\mathbf{x}_{i},\mathbf{x}_{l})\leq d_{max}^{C_{j}}$, where $d_{max}^{C_{j}}$ is the largest distance in the cluster’s minimum spanning tree (MST). The resulting graph accurately describes the connectivity within the cluster at the density where the last point joins the cluster. The space tree constructed by HDBSCAN* is used to retrieve these edges efficiently. The core approximation graph adds all edges with $d_{mreach}(\mathbf{x}_{i},\mathbf{x}_{j})\leq\max{\{\kappa{(\mathbf{x}_{i})},\kappa{(\mathbf{x}_{j})}\}}$. The resulting graph accurately describes the connectivity in the cluster’s MST. Alternatively, this graph can be considered as the cluster’s subgraph from the $k$-nearest neighbour graph over the entire data set. HDBSCAN* already extracted these edges when the core distances were computed, so this approach has a lower additional cost.
We can now recover the eccentricity contour tree as if it is were density contour tree by applying HDBSCAN*’s clustering steps to the cluster approximation graph with its edges weighted by $\min{\{e(\mathbf{x}_{i}),e(\mathbf{x}_{l})\}}$. This weighting ensures an edge has the eccentricity of the least eccentric point it connects. Specifically, we adapt McInnes and Healy (2017)’s Union-Find data structure to construct a single linkage dendrogram. The resulting hierarchy is simplified using a minimum branch size $m_{b}$ to recover the condensed branching hierarchy shown in Fig. 2(d).
@@
HDBSCAN*’s EOM and leaf strategies are used to compute branch labels and membership probabilities from these condensed hierarchies. Points that enter the filtration after the selected branches have connected—i.e. points with the noise label—are given a single non-noise label representing the cluster’s centre. Finally, the cluster and branch labels are combined. By default, points in clusters with two or fewer branches are given a single label because two branches are expected in all clusters, indicating the outsides growing towards each other. The label sides as branches parameter can be used to turn off this behaviour and separate the ends of elongated clusters in the labelling. The cluster and branch probabilities are combined by taking their average value (Fig. 3(a)).
@@
Other labelling and probability combinations are possible. For example, the cluster and branch probability product more strongly emphasises the outsides of the branches (Fig. 3(b)). As in McInnes et al. (2017), FLASC supports computing branch membership vectors that describe how strongly a point $\mathbf{x}_{i}\in C_{j}$ belongs to each branch $B_{b}\subset C_{j}$. These membership values are based on the geodesic distances in the cluster approximation graph $\mathbf{G}_{k}^{C_{j}}$: $d_{geo}(\mathbf{r}_{B_{b}},\mathbf{x}_{i})$, where $\mathbf{r}_{B_{b}}$ is the branch’ root, i.e., the point closest to the branch’s membership-weighted average $\bar{\mathbf{x}}_{B_{b}}$. The branch membership vectors can be used to label central points by the closest branch root, as in Fig. 3(c). Alternatively, a softmax function can be used to convert $d_{geo}(\mathbf{r}_{B_{b}},\mathbf{x}_{i})$ into the membership probabilities:
$p(\mathbf{x}_{i},B_{b})=\frac{e^{c_{b}(\mathbf{x}_{i},B_{b})/t}}{\sum_{B_{l}\in C_{j}}{e^{c_{b}(\mathbf{x}_{i},B_{l})/t}}},$ (4)
where $c_{b}(\mathbf{x}_{i},B_{b})=1/d_{geo}(\mathbf{r}_{B_{b}},\mathbf{x}_{i})$ and $t$ is a temperature parameter (Fig. 3(d)).
@@
Low persistent branches can be ignored using a branch selection persistence parameter, analogous to HDBSCAN*’s cluster selection epsilon (Malzer and Baum, 2020). As branches do not necessarily start at zero centrality, branch selection persistence describes the minimum eccentricity range, rather than a single eccentricity threshold value. The procedure that applies the threshold simplifies the condensed branch hierarchy until all leaves have a persistence larger than the threshold.
Stability
(a) Probability average
(b) Probability product
(c) Geodesic labels
(d) Geodesic membership
@@
Figure 3: Different ways to combine cluster and branch membership probabilities. The cluster and branch probability average (a) and product (b) are visualised with desaturation. (c) Points labelled by the geodesically closest branch root—i.e., the point closest to the branch’s weighted average—and desaturated as in (a). (d) Weighted branch membership for the orange branch is visualised by transparency. Branch memberships are computed from the traversal distance to the branch’s root.
@@
Stability is an important property of algorithms indicating that their output differs only slightly when the input changes slightly. Two notions of stability are relevant for FLASC: (1) the algorithm has to provide similar results when run repeatedly on (different) samples of an underlying distribution, and (2) the detected branch hierarchies have to represent the clusters’ underlying topology accurately. The deterministic density-based design of FLASC provides stability in the first sense.
@@
Vandaele et al. (2021) analysed the second notion of stability for graph-based branch detection, explaining that the graph approximation should accurately represent the underlying shape and the graph-based centrality function should accurately describe the points’ centrality in a cluster’s metric space $(C_{j},d_{mreach})$. For the normalised centrality used by Carlsson (2014), Vandaele et al. (2021) show that the bound on the bottleneck distances between true and empirical persistence diagrams is tight if the metric distortion induced by the graph and its maximum edge weight is small.
@@
The eccentricity function (Equation (3)) is more complex to analyse. It is a $\mathrm{c}$-Lipschitz-continuous function when considered over the cluster centrality graph’s edges:
$|\max{\{e(\mathbf{x}_{i}),e(\mathbf{x}_{j})\}}-\max{\{e(\mathbf{x}_{k}),e(\mathbf{x}_{l})\}}|\leq\mathrm{c}\;d_{mreach}(\mathbf{x}_{i},\mathbf{x}_{l}),$
@@
where $\mathrm{c}$ is a constant describing the continuity, $(\mathbf{x}_{i},\mathbf{x}_{j})\in\mathbf{G}_{k}^{C_{j}}$ and $(\mathbf{x}_{k},\mathbf{x}_{l})\in\mathbf{G}_{k}^{C_{j}}$, and the mutual reachability between $\mathbf{x}_{i}$ and $\mathbf{x}_{l}$ is the largest of the four points. However, the position of the cluster’s centroid $\bar{\mathbf{x}}_{C_{j}}$ influences how well the function represents the centrality in the cluster’s metric space $(C_{j},d_{mreach})$. We aim to show the current approach strikes a good balance between computational cost and stability in the experiments presented in the next section.
Computational Complexity
@@
The algorithm’s most computationally expensive steps are constructing the full cluster approximation graphs and computing their single linkage hierarchies. Naively, the worst-case complexity for creating a cluster approximation graph is $\mathcal{O}(n_{c}^{2})$, where $n_{c}$ is the number of points in the cluster. Usually, the average case is much better because the approximation graphs rarely are fully connected. After all, HDBSCAN*’s noise classification limits the density range within the clusters. Furthermore, the space tree that we re-use from the HDBSCAN* clustering step provides fast asymptotic performance for finding the graph’s edges. The exact run-time bounds depend on the data properties. They are challenging to describe (as explained in McInnes and Healy (2017)), but an average complexity proportional to $\mathcal{O}(n_{e}\log{N})$ is expected, where $n_{e}$ is the number of edges in the approximation graph. Computing single linkage hierarchies from the cluster centrality graphs is possible in $\mathcal{O}(n_{e}\alpha(n_{e}))$ using McInnes et al. (2017)’s Union-Find implementation adapted to ignore edges between data points that are already in the same connected component ($\alpha$ is the inverse Ackermann function). Like McInnes and Healy (2017), we feel confident that FLASC achieves sub-quadratic complexity on average, which we demonstrate in a practical example in the next section.
Experiments
Stability and computational performance are essential properties for clustering algorithms to be useful in practice. This section presents two synthetic benchmarks that compare these properties between FLASC’s, HDBSCAN*, and kMeans. In addition, we demonstrate two exploration use cases where detecting branch-based subgroups and branch hierarchies helps to understand the data’s structure.
FLASC Stability
This first synthetic benchmark compares the ability of FLASC, HDBSCAN*, and kMeans to detect branch-based subgroups that do not contain a density maximum. The benchmark is designed to answer the following research question:
How does FLASC’s branch detection ability compare to kMeans and HDBSCAN* in terms of accuracy and stability?
kMeans was selected for this comparison because it is a popular, fast clustering algorithm that should be able to detect branch-based subgroups when given the correct number of subgroups a priori. HDBSCAN* was included to show that it cannot reliably find subgroups that do not contain a density maximum, demonstrating that our branch-detection post-processing step enables the algorithm to find an additional type of pattern.
We expect both kMeans and FLASC to be able to detect branching structures, while HDBSCAN* will struggle due to the lack of density maxima in the to-be-detected subgroups. Furthermore, we expect FLASC to have a higher stability than kMeans—in terms of difference between predictions on multiple samples of the same underlying distribution—as it does not rely on converging from a random initialisation. The next two subsections explain how the data for this benchmark was generated and how we measured the algorithm’s accuracy and stability, respectively.
Datasets
@@
Two-dimensional data sets containing a single cluster with $3$ branches laid out as a three-pointed star were generated (Fig. 4). The branches span from the centre outwards and are spread out equally across the $2$D plane—i.e., the angle between adjacent branches is $120$ degrees. Each branch has a length $b_{l}$ and contains $100$ points exponentially spaced from the inside out. Consequently, the density is highest at the cluster’s centre and lowest in the branch ends, and this difference increases with the branch length. Normally distributed noise ($\mu=0$) is added to the points’ coordinates using a noise ratio parameter $n_{r}$ to determine the distribution’s standard deviation $\sigma$:
$n_{r}=\frac{4\sigma}{\sqrt{\frac{1}{2}}b_{l}}.$ (5)
@@
Here, $n_{r}$ indicates that approximately $95\%$ of the sampled noise values fall within ±$\pm\sqrt{\frac{1}{2}}n_{r}b_{l}$ of zero. Consequently, approximately $90\%$ of points are moved less than $n_{r}b_{l}$ from their original 2D position.
Evaluation and settings
Figure 4: Explanatory figure for the centroid spread stability measure. Predicted centroids—weighted average coordinates—are computed for each predicted subgroup and assigned to the closest ground-truth centroid. The 95 percentile centroid spread can be interpreted by translating each group’s average to the origin and finding the radius for a circle that includes 95% of the predicted centroids.
@@
The data sets were sampled with varying branch lengths $b_{l}$ ($2$ to $100$ in $10$ exponentially spaced steps) and noise ratios $n_{r}$ (00 to $1$ in $10$ exponentially spaced steps). Ten data sets were sampled for each parameter combination, resulting in $1000$ point clouds. FLASC was tuned to find a single cluster by enabling the allow single cluster parameter and setting min samples $k=5$ and min cluster size $m_{c}=25$. Its branch detection parameters were optimised in a grid search: using the full and core approximation graphs, both branch selection strategies, and varying the min branch size $m_{b}$ between $2$ to $24$ in steps of $2$. Labels for central points were set to the geodesically closest branch root (as in Fig. 3(c)). HDBSCAN* was tuned to find multiple clusters using min samples $k=5$. Its other parameters were optimised in a grid search: varying min cluster size $m_{c}=$ between $2$ to $24$ in steps of $2$, and using both cluster selection strategies. For, kMeans, we set $k=3$, using our prior knowledge that there are three branch-based subgroups in the data.
@@
For each evaluation, the resulting data point labels and probabilities were stored and used to computed the Adjusted Rand Index (ARI) (Hubert and Arabie, 1985). ARI values describe the agreement between ground truth and assigned labels adjusted for chance. In addition, the algorithm’s stability was measured using the spread of the predicted branch centroids (see, Fig. 4). Here, the weighted average coordinates per detected subgroup were used as predicted centroid. These predicted centroids were assigned to the closest ground-truth centroid. Their spread was quantified as the distance between the predicted centroids and their unweighted average per ground-truth centroid. This value was normalised by the branch length $b_{l}$ to allow comparisons across branch lengths.